# Particle Spectra in the Integrated Hydrokinetic Model at RHIC BES Energies

### v1

@@ Abstract
This paper focuses on the description and prediction of particle production in Au+Au collisions at various center-of-mass energies ($\sqrt{s_{\text{NN}}}=7.7,11.5,14.5,19.6,27.0,\text{ and }39.0$ GeV), using the Integrated HydroKinetic model (iHKM). The iHKM, initially created for description of ultratelativistic nucleus-nucleus collisions, has been extended to cover intermediate and low relativistic energies, ranging from 2.5 to 50 GeV per colliding nucleon pair, and in cases where this distinction is needed, we refer to the modified version as iHKMe. The study explores the sensitivity of model predictions to variations in key parameters, including the relaxation time and thermalization degree during the prethermal stage, as well as the nature of the equation of state transition in the hydrodynamic phase. The iHKM successfully reproduces experimental results from the STAR experiment within the RHIC Beam Energy Scan program, with particular emphasis on midrapidity ($|y|<0.1$) observables such as particle multiplicity densities ($dN/dy$), average transverse momenta ($\langle p_{T}\rangle$), and particle yield ratios. In addition, the paper examines the evolution of chemical and kinetic freeze-out across different collision energies and centrality classes. These results offer a consistent description of the collective properties of strongly interacting matter formed in heavy-ion collisions, spanning a wide range of energies and baryon chemical potentials at RHIC.
@@ I Introduction
It is widely considered that relativistic heavy-ion collisions, such as those conducted at the Large Hadron Collider (LHC) and the Relativistic Heavy Ion Collider (RHIC), create new forms of strongly interacting matter. Under extreme temperatures and densities, deconfined quarks and gluons can form a nearly equilibrated state known as the quark-gluon plasma (QGP) [1]. In contrast, under normal Earth conditions, quarks and gluons are confined within color-neutral hadrons due to the phenomenon of confinement.
In the quantum chromodynamics (QCD) phase diagram, the QGP and the hadron-resonance gas (HRG) phases are connected via a smooth crossover transition at vanishing baryon chemical potential [2], a scenario realized in collisions at LHC energies and the top RHIC energies. However, at higher baryon densities, the system is expected to undergo a genuine phase transition, potentially of first order [3, 4]. If such a phase transition line exists in the QCD phase diagram, it must end at a critical point, beyond which the transition becomes a crossover. Ongoing and upcoming experiments at lower collision energies, typically in the range of several GeV per nucleon pair, aim to identify clear signals of this phase transition and locate the critical endpoint.
With the launch of the Beam Energy Scan (BES) program at RHIC, three primary objectives were outlined [5]. The first goal is to explore the QCD phase diagram by varying the center-of-mass collision energy, $\sqrt{s_{NN}}$, in the range from 7.7 to 39 GeV. This variation enables access to different regions of temperature $T$ and baryon chemical potentials $\mu_{B}$ during the system’s evolution, facilitating the search for the onset of deconfinement. The second goal is to investigate fluctuations and enhanced susceptibilities that are expected to increase near the QCD critical point [6, 7, 8]. However, such critical signatures may be suppressed due to conservation laws and the finite size and lifetime of the system. This challenge leads to the third goal: to identify observables sensitive to the softening of the QCD equation of state (EoS), which may signal the presence of a first-order phase transition.
@@
Relativistic hydrodynamics has emerged as the most effective framework for describing the collective expansion of the strongly interacting matter created in these collisions. However, it comes with notable limitations: it is primarily applicable to soft observables (low transverse momentum, $p_{T}$) and assumes local thermal equilibrium. To address these constraints, theorists have developed hybrid models that integrate relativistic hydrodynamics, used for the hot, dense equilibrated stage, with microscopic transport models that describe out-of-equilibrium dynamics. Such non-equilibrium conditions are expected both in the very early pre-equilibrium stage, before thermalization occurs, and in the late afterburner stage, when the system becomes dilute and departs from equilibrium due to expansion.
@@
At the LHC and the highest RHIC energies, the standard model of heavy-ion collisions is well established and widely accepted [9]. This model typically divides the collision process into several distinct stages: the formation of a non-equilibrium state during the penetration of the colliding nuclei, subsequent thermalization process leading to formation of a quark-gluon plasma (QGP), hydrodynamic expansion of the resulting fireball, hadronization of the QGP into a hadron resonance gas (HRG), and finally, the hadronic cascade stage [10, 11, 12, 13, 14, 15]. As the collision energy decreases, this overall picture remains qualitatively similar but becomes significantly more complex. First, the characteristic time for two gold nuclei to fully overlap is approximately $10^{-3}$ fm/$c$ at $\sqrt{s_{NN}}=5.02$ TeV, while at $\sqrt{s_{NN}}=14.5$ GeV at BES RHIC it exceeds 1.5 fm/$c$. As a result, the system becomes highly inhomogeneous, not only in the transverse but also in the longitudinal direction. Second, the lower energy density leads to a slower thermalization process, which may significantly shorten, or even eliminate the hydrodynamic phase. At collision energies of just a few GeV, it is plausible that the pre-equilibrium stage dominates the evolution, preventing the system from ever reaching full thermal equilibrium.
@@
In a recent work [16], we proposed a novel framework to describe the full evolution of strongly interacting matter in relativistic heavy-ion collisions at intermediate and low energies, starting from the initial collision to the decoupling of free-streaming hadrons. In this extended version of the integrated hydrokinetic model, referred to as iHKMe, we adopt standard hydrodynamics to model the bulk evolution while simultaneously incorporating the UrQMD model to account for non-equilibrium dynamics. Unlike many hybrid approaches [17, 18, 19, 20], which segment the evolution into distinct, sequential phases, iHKMe integrates microscopic and macroscopic descriptions concurrently across both equilibrium and non-equilibrium stages. In our approach, the system initially exists far from equilibrium and is gradually thermalized, transitioning into a hydrodynamic regime [^1] Another approach to continuous transition from UrQMD to hydrodynamics is discussed in [21]. Since the thermalization rate is not well constrained, it is treated as a free parameter in the model. By tuning this parameter, iHKMe enables a smooth interpolation between scenarios dominated by rapid thermalization (as at LHC energies) and those that remain largely non-equilibrated throughout the evolution, typical of low-energy collisions.
@@
This study serves as a continuation of our previous work [16], in which we introduced an extended version of the integrated hydrokinetic model. In the present paper, we focus on describing the transverse momentum spectra of particles produced in Au-Au collisions at $\sqrt{s_{NN}}=7.7-39$ GeV, as part of the RHIC Beam Energy Scan (BES) program. We calibrate the model’s free parameters using two different equations of state: one featuring a smooth crossover transition [22], and the other incorporating a first-order phase transition with a soft equation of state [23].
@@ II Model
The iHKM model was originally developed to study ultrarelativistic heavy-ion collisions [15]. We have recently extended the model to accommodate intermediate and low collision energies, broadening its applicability down to a few GeV. A detailed description of the extended model can be found in [16]. In this paper, we present a brief overview of the model’s five evolutionary stages, discuss the possible overlap of these stages at lower collision energies, and introduce the key free parameters that control the model’s dynamics.
@@ II.0.1 Initial pre-equilibrium dynamics
One of the key differences between ultrarelativistic heavy-ion collisions at the LHC and the lower-energy experiments of the RHIC BES program lies in the very different relative velocities of colliding nuclei (and correspondingly space-time Lorentz factors for them), that, according to QCD approaches, leads to different physical pictures of nucleus-nucleus collisions at different energies. First, it concerns the different initial states for the colliding process. At TeV-scale energies, partons from the two nuclei interact only during the extremely short overlap time, approximately $10^{-3}$ fm/$c$, which is negligible compared to the full evolution timescale of the system (typically exceeding 10 fm/$c$). In contrast, at lower energies, the scenario becomes significantly more complex. The duration of the nuclear overlap, which increases with decreasing beam energy, can be estimated using the following equation:
@@
where $R_{N}$ is the radius of a nucleus and $m_{N}$ is the nucleon mass. A straightforward calculation shows that at intermediate energies, around $\sqrt{s_{NN}}\sim 10$ GeV, the overlap time already exceeds $1.5$ fm/$c$. Thermalization can only begin during this stage, meaning that a purely hydrodynamic description is not yet applicable. Consequently, a microscopic approach must be employed to model the early, out-of-equilibrium dynamics of the system.
In the iHKM for ultra-relativistic energies, the initial state of matter is generated using the Monte Carlo Glauber model, implemented in the GLISSANDO II package [24], featuring a longitudinally boost-invariant distribution in coordinate space and a highly anisotropic momentum distribution inspired by the Color Glass Condensate framework [25]. In the current version, event-by-event UrQMD simulations are employed to describe the initial non-equilibrium dynamics up to a proper time $\tau_{0}$, which marks the onset of the thermalization stage, discussed in detail later.
The output of UrQMD simulations is used to construct a stress-energy tensor and baryon charge current, which are then used as initial conditions in the hydrodynamic equations. Summing over all particles (labeled with index $i$) within each proper time step $\Delta\tau$,
$\left|\sqrt{t_{i}^{2}-z_{i}^{2}}-\tau\right|<\Delta\tau/2,$ (2)
we obtain
$T^{\mu\nu}_{\text{urqmd}}(\tau;x_{j})$ $=\sum_{i}\frac{p_{i}^{\mu}p_{i}^{\nu}}{p_{i}^{0}}{\cal K}_{ij},$ (3)
$J^{\mu}_{\text{urqmd}}(\tau;x_{j})$ $=\sum_{i}B_{i}\frac{p_{i}^{\mu}}{p_{i}^{0}}{\cal K}_{ij}.$ (4)
Here, $T^{\mu\nu}_{\text{urqmd}}$ represents the stress-energy tensor at proper time step $\tau_{j}$, and $J^{\mu}_{\text{urqmd}}(\tau;x_{j})$ is the baryon current  [^2] In this paper, we do not consider separate equations for other conserved charges. The electric charge density is assumed to be proportional to the baryon density, $n_{q}=\frac{Z}{A}n_{b}$, where $\frac{Z}{A}=0.4$ for gold nuclei. The strangeness is always locally zero.. Here, $p^{\mu}_{i}$ denotes the four-momentum of particle $i$, and $B_{i}$ represents its baryon charge. We also utilize a ${\cal K}_{ij}$ for particle smearing to generate a relatively smooth or averaged over the ensemble of similar events tensors.
@@
. In this expression, $r_{ij}^{\mu}=x^{\mu}_{i}-x^{\mu}_{j}$ represents the radius vector between the particle and the center of the cell ($x_{i},x_{j}\in\sigma^{\mu}$), $u_{i\lambda}$ is four-velocity of particle, and $n_{j}^{\lambda}$ is a normal vector to the hypersurface on the lattice $\Delta\sigma^{\mu}$
@@
$n^{\mu}$ $=\left(\tau\cosh\eta,0,0,\tau\sinh\eta\right),$ (7)
In Eq. (5), a free scalar parameter of the model $R$ is introduced. It is responsible for the smoothness of the initial collisions. Lastly, let us specify that $\sum_{j}K_{ij}=1$, ensuring no noticeable violations of conservation laws in this procedure.
@@ II.0.2 Thermalzaition
During the thermalization stage, the system is assumed to consist of two coexisting phases. The first is a non-equilibrium component that evolves according to the UrQMD model and is described by Eq. (3). The second represents near-equilibrium matter governed by relativistic hydrodynamics. Its tensors can be written as follows
$T^{\mu\nu}_{\text{hydro}}(x)=\left(\epsilon+p\right)u^{\mu}u^{\nu}-pg^{\mu\nu}+\pi^{\mu\nu}.$ (8)
@@
where $\epsilon$, $p$, and $n_{B}$ represent the local energy density, pressure, and baryonic density, respectively. $u^{\mu}$ denotes the four-velocity of the fluid, $g^{\mu\nu}$ is the metric tensor, and $\pi^{\mu\nu}$ stands for the shear-stress tensor. The energy density, baryon density, and pressure are connected through the equation of state:
$p=p(\epsilon,n_{B}).$ (10)
@@
where ${\cal P}_{\tau}={\cal P}(\tau)$ is a weight function such that ${\cal P}(\tau_{0})=1$ at the start of the thermalization stage, ${\cal P}(\tau_{th})=0$ at the end, and $10<{\cal P}(\tau_{0}<\tau<\tau_{th})<$ 1 in between. So this weight function regulates how matter is pumped from one component to another during the thermalization. As in our previous papers [26, 27] we utilize an ansatz for ${\cal P}(\tau)$ inspired by the Boltzmann equation in relaxation time approximation.
${\cal P}(\tau)=\left(\frac{\tau_{th}-\tau}{\tau_{th}-\tau_{0}}\right)^{\frac{\tau_{th}-\tau_{0}}{\tau_{rel}}}.$ (13)
This function involves three free parameters: the onset of thermalization $\tau_{0}$, end of thermalization $\tau_{th}$, and its speed encoded in relaxation time $\tau_{rel}$.

### v2

@@ Abstract
We study light-hadron production in Au+Au collisions at $\sqrt{s_{NN}}=7.7$–39 GeV using an extended Integrated HydroKinetic Model (iHKMe). Focusing on transverse momentum spectra, we investigate the sensitivity to key model parameters, in particular the thermalization time scale. We consider two distinct equations of state: one featuring a crossover and the other a first-order phase transition. In both cases, thermalization begins shortly before full nuclear overlap and lasts approximately 1 fm/$c$ across all energies. Both equations of state provide a similarly good description of the soft particle momentum spectra once the other parameters are slightly adjusted. The most pronounced differences arise at the lower RHIC BES energy of $\sqrt{s_{NN}}=7.7$ GeV, particularly in proton and kaon yields, reflecting their sensitivity to the freeze-out parameters.
@@ I Introduction
Relativistic heavy-ion collisions, such as those at the Large Hadron Collider (LHC) and the Relativistic Heavy Ion Collider (RHIC), are known to create new forms of strongly interacting matter. At extreme temperatures and densities, quasi-free quarks and gluons can form a nearly equilibrated state known as the quark–gluon plasma (QGP) [1]. In contrast, under ordinary terrestrial conditions, quarks and gluons remain confined within color-neutral hadrons [2, 3].
In the quantum chromodynamics (QCD) phase diagram, lattice QCD calculations indicate that the quark–gluon plasma and hadron–resonance gas (HRG) phases are connected by a smooth crossover transition at vanishing baryon chemical potential [3, 4, 5], a situation realized at LHC and top RHIC energies. However, at higher baryon densities, the system is expected to undergo a genuine phase transition, potentially of first order [6, 7, 8]. If such a phase transition line exists in the QCD phase diagram, it must end at a critical point, beyond which the transition becomes a crossover. Ongoing and upcoming experiments [9, 10, 11] at lower collision energies, typically in the range of several GeV per nucleon pair, aim to identify clear signals of this phase transition and locate the critical endpoint.
With the launch of the Beam Energy Scan (BES) program at RHIC, three primary objectives were outlined [12, 13]. The first goal is to explore the QCD phase diagram by varying the center-of-mass collision energy, $\sqrt{s_{NN}}$, in the range from 7.7 to 39 GeV. This variation enables access to different regions of temperature $T$ and baryon chemical potentials $\mu_{B}$ during the system’s evolution, facilitating the search for the onset of deconfinement. The second goal is to investigate fluctuations and enhanced susceptibilities that are expected to increase near the QCD critical point [14, 15, 16]. However, such critical signatures may be suppressed due to conservation laws and the finite size and lifetime of the system. This challenge leads to the third goal: to identify observables sensitive to the softening of the QCD equation of state (EoS), which may signal the presence of a first-order phase transition.
@@
Relativistic hydrodynamics has emerged as the most effective framework for describing the collective expansion of the strongly interacting matter created in these collisions. However, it comes with notable limitations: it is primarily applicable to soft observables (low transverse momentum, $p_{T}$) and assumes local thermal equilibrium [17]. To address these constraints, theorists have developed hybrid models that integrate relativistic hydrodynamics, used for the hot, dense equilibrated stage, with microscopic transport models that describe out-of-equilibrium dynamics. Such non-equilibrium conditions are expected both in the very early pre-equilibrium stage, before thermalization occurs, and in the late afterburner stage, when the system becomes dilute and departs from equilibrium due to expansion.
@@
At the LHC and the highest RHIC energies, the standard model of heavy-ion collisions is well established and widely accepted [18, 19, 20, 21]. This model typically divides the collision process into several distinct stages: the formation of a non-equilibrium state during the penetration of the colliding nuclei, subsequent thermalization process leading to formation of QGP, hydrodynamic expansion of the resulting fireball, hadronization of the QGP into HRG, and finally, the hadronic cascade stage [22, 23, 24, 25, 26, 21, 27, 28, 29]. As the collision energy decreases, this overall picture remains qualitatively similar but becomes significantly more complex. First, the characteristic time for two gold nuclei to fully overlap is approximately $10^{-3}$ fm/$c$ at $\sqrt{s_{NN}}=5.02$ TeV, while at $\sqrt{s_{NN}}=14.5$ GeV at BES RHIC it exceeds 1.5 fm/$c$. Consequently, the system becomes highly inhomogeneous not only in the transverse but also in the longitudinal direction, leading to a significant violation of boost invariance (i.e. a breakdown of the Bjorken picture [30]). Second, the lower energy density leads to a slower thermalization process, which may significantly shorten, or even eliminate, the hydrodynamic phase. At collision energies of just a few GeV, it is plausible that the pre-equilibrium stage dominates the evolution, preventing the system from ever reaching full thermal equilibrium.
@@
In a recent study [31], we introduced an extended framework to describe the complete space–time evolution of the system created in relativistic heavy-ion collisions at intermediate and low energies, from the initial nuclear overlap to the final decoupling of free-streaming hadrons. This extended version of the integrated hydrokinetic model, denoted iHKMe, employs relativistic hydrodynamics for the bulk equilibrated medium evolution, while the UrQMD cascade [32] is utilized to account for the essential non-equilibrium dynamics. Unlike many hybrid approaches [33, 34, 35, 36], which segment the evolution into distinct phases in space-time, iHKMe integrates microscopic and macroscopic descriptions concurrently. In our approach, the system initially exists far from equilibrium and gradually thermalizes, transitioning into a hydrodynamic regime [^1] Other approaches to a continuous transition from UrQMD to hydrodynamics are discussed in [35, 37]. Since the thermalization rate is not well constrained, it is treated as a free parameter in the model. By tuning this parameter, iHKMe enables a smooth interpolation between scenarios dominated by rapid thermalization (as at LHC energies) and those that remain largely non-equilibrated throughout the evolution, as expected for very low-energy collisions.
@@
This study continues our previous work [31], in which we introduced an extended version of the integrated hydrokinetic model. In the present paper, we focus on describing the transverse momentum spectra of particles produced in Au-Au collisions at $\sqrt{s_{NN}}=7.7-39$ GeV, as part of the RHIC Beam Energy Scan program. We calibrate the model’s free parameters using two different equations of state: one featuring a smooth crossover transition [38], and the other incorporating a first-order phase transition with a soft equation of state [39].
@@ II Model
The iHKM model was originally developed to study ultrarelativistic heavy-ion collisions [29]. We have recently extended the model to accommodate intermediate and low collision energies, broadening its applicability down to a few GeV. A detailed description of the extended model can be found in [31]. In this paper, we provide only a brief overview of the model’s five evolutionary stages, discuss the potential overlap of these stages at lower collision energies, and introduce the key free parameters that govern the model’s dynamics.
@@ II.0.1 Initial pre-equilibrium dynamics
One of the key differences between ultrarelativistic heavy-ion collisions at the LHC and the lower-energy experiments of the RHIC BES program lies in the very different relative velocities of colliding nuclei (and correspondingly space-time Lorentz factors for them), which, according to QCD approaches, leads to different physical pictures of nucleus-nucleus collisions at different energies. First, it concerns the different initial states for the colliding process. At TeV-scale energies, partons from the two nuclei interact only during the extremely short overlap time, approximately $10^{-3}$ fm/$c$, which is negligible compared to the full evolution timescale of the system (typically exceeding 10 fm/$c$). In contrast, at lower energies, the scenario becomes significantly more complex. The duration of the nuclear overlap, which increases with decreasing beam energy, can be estimated using the following equation:
@@
where $R_{N}$ is the radius of a nucleus and $m_{N}$ is the nucleon mass. Here we adopt the proper time $\tau$ and space-time rapidity $\eta$ variables, following the conventional Bjorken coordinate system for longitudinally expanding systems [30]. A straightforward calculation shows that at intermediate energies, around $\sqrt{s_{NN}}\sim 10$ GeV, the overlap time already exceeds $1.5$ fm/$c$. Thermalization only begins during this stage, meaning that a purely hydrodynamic description is not yet applicable. Consequently, a microscopic approach must be employed to model the system’s early, out-of-equilibrium dynamics.
In the iHKM for ultra-relativistic energies, the initial state of matter is generated using the Monte Carlo Glauber model, implemented in the GLISSANDO II package [40], featuring a longitudinally boost-invariant distribution in coordinate space and a highly anisotropic momentum distribution inspired by the Color Glass Condensate framework [41, 42, 43, 44]. In the current version, event-by-event UrQMD cascade simulations are employed to describe the initial non-equilibrium dynamics up to a proper time $\tau_{0}$, which marks the onset of the thermalization stage, discussed in detail later.
The output of UrQMD simulations is used to construct the stress-energy tensor and baryon charge current, which serve as initial conditions for the hydrodynamic equations [24, 45, 46]. At each step of proper time $\tau_{j}\pm\Delta\tau/2$, we select all particle tracks [^2] Here we refer to the particle positions at each step of Cartesian time $t$ obtained from UrQMD. For consistency, the same time step is used in UrQMD and in the subsequent hydrodynamic simulations, $\Delta t=\Delta\tau$. (labeled by index $i$) that satisfy
$\left|\sqrt{t_{i}^{2}-z_{i}^{2}}-\tau_{j}\right|<\frac{\Delta\tau}{2}\,.$ (2)
The space-time distributions of the macroscopic fields are then constructed as
$T^{\mu\nu}_{\text{urqmd}}(x_{j})$ $=\sum_{i}\frac{p_{i}^{\mu}p_{i}^{\nu}}{p_{i}^{0}}\,{\cal K}_{ij},$ (3)
$J^{\mu}_{\text{urqmd}}(x_{j})$ $=\sum_{i}B_{i}\frac{p_{i}^{\mu}}{p_{i}^{0}}\,{\cal K}_{ij}.$ (4)
Here, $T^{\mu\nu}_{\text{urqmd}}$ represents the stress-energy tensor at the space-time point $x^{\mu}_{j}$, $J^{\mu}_{\text{urqmd}}(x_{j})$ is the baryon current [^3] In this paper, we do not consider separate equations for other conserved charges. The electric charge density is assumed to be proportional to the baryon density, $n_{q}=\frac{Z}{A}n_{b}$, where $\frac{Z}{A}\approx 0.4$ for gold nuclei. The strangeness is always locally zero. , $p^{\mu}_{i}$ denotes the four-momentum of particle $i$, and $B_{i}$ represents its baryon charge. We also utilize a ${\cal K}_{ij}$ for particle smearing to generate a relatively smooth or averaged over the ensemble of similar events tensors.
@@
In this expression, $r_{ij}^{\mu}=x^{\mu}_{i}-x^{\mu}_{j}$ represents the four-vector connecting the particle and the center of the cell ($x_{i},x_{j}\in\sigma^{\mu}$) in $3+1$-dimensional spacetime, $u_{i\lambda}$ is four-velocity of particle, and $n_{j}^{\lambda}$ is a normal vector to the hypersurface on the lattice $\Delta\sigma^{\mu}$
@@
$n^{\mu}$ $=\left(\tau\cosh\eta,0,0,\tau\sinh\eta\right)\,.$ (7)
In Eq. (5), a free scalar parameter of the model $R$ is introduced. It is responsible for the smoothness of the initial collisions. Lastly, let us specify that $\sum_{j}K_{ij}\approx 1$, ensuring no noticeable violations of conservation laws in this procedure [31].
@@ II.0.2 Thermalzaition
During the thermalization stage, the system is assumed to consist of two coexisting phases. The first is a non-equilibrium component that evolves according to the UrQMD model, as described by Eq. (3). The second corresponds to near-equilibrium matter, which is governed by relativistic hydrodynamics. Its macroscopic tensors are given by
$T^{\mu\nu}_{\text{hydro}}(x)=(\epsilon+p)u^{\mu}u^{\nu}-pg^{\mu\nu}+\pi^{\mu\nu},$ (8)
@@
where $\epsilon$, $p$, and $n_{B}$ denote the local energy density, pressure, and baryon density, respectively; $u^{\mu}$ is the fluid four-velocity, $g^{\mu\nu}$ is the metric tensor, and $\pi^{\mu\nu}$ is the shear-stress tensor. In this work, we neglect other dissipative effects such as bulk viscosity, heat flow, and baryon diffusion, focusing solely on the shear-stress contribution in the hydrodynamic evolution. The pressure and energy density are related through the equation of state.
@@
where ${\cal P}_{\tau}={\cal P}(\tau)$ is a weight function such that ${\cal P}(\tau_{0})=1$ at the start of the thermalization stage, ${\cal P}(\tau_{\text{th}})=0$ at the end, and $0<{\cal P}(\tau_{0}<\tau<\tau_{\text{th}})<1$ in between. So this weight function regulates how matter is pumped from one component to another during the thermalization. As in our previous papers [47, 48] we utilize an ansatz for ${\cal P}(\tau)$ inspired by the Boltzmann equation in relaxation time approximation.
${\cal P}(\tau)=\left(\frac{\tau_{\text{th}}-\tau}{\tau_{\text{th}}-\tau_{0}}\right)^{\frac{\tau_{\text{th}}-\tau_{0}}{\tau_{\text{rel}}}}.$ (12)
This function involves three free parameters: the onset of thermalization $\tau_{0}$, the end of thermalization $\tau_{\text{th}}$, and its rate encoded in the relaxation time $\tau_{\text{rel}}$.